OpenStudy (anonymous):

Suppose that at the end of any given year an automobile is only worth 80% of it's value a the beginning of that year. If a car cost 21,000 new, what is it's value at the end of five years?

OpenStudy (dumbcow):

21000*(1-0.8)^5

OpenStudy (dumbcow):

ignore that post sorry

OpenStudy (anonymous):

lol.

OpenStudy (dumbcow):

21000*(0.8^5) :)

OpenStudy (radar):

1st yr .8x$21000=$16800 2nd yr .8 x 16800=13440 3rd yr .8 x 13440=10752 4th yr .8 x 10752=8601.60 5th yr .8 x 8601.60=$6881.28

OpenStudy (anonymous):

hmmmmm...........

OpenStudy (radar):

Or as dumbcow stated\[21000*.8^{5}\]

OpenStudy (anonymous):

BAM! both of you two were right! I just had to round down. Thanks so much on this one guys!

OpenStudy (anonymous):

A company offers a starting yearly salary of 31,000, with raises of 2,000 per year. Find the total salary over a ten-year period.

OpenStudy (anonymous):

Something tricky about this one, it's like 300k, or 400k sumthin...

OpenStudy (dumbcow):

First find salary at end of ten-year period A10 = 31000+(10-1)*2000 A10 = 49000 Then use series sum formula Sum = 1/2*n*(A1+A10) Sum = 1/2*10*(31000+49000) ...

OpenStudy (anonymous):

.......

OpenStudy (dumbcow):

thats left for you to compute ;)

OpenStudy (anonymous):

I am a bottom line kind of girl..... :)

OpenStudy (anonymous):

It doesn't add up.......hmmm...

OpenStudy (anonymous):

Sigh....

OpenStudy (dumbcow):

what did you come up with?

OpenStudy (anonymous):

Hello.

OpenStudy (anonymous):

Ah well.......I came up with a bunch of letters I don't understand.

OpenStudy (dumbcow):

what do you mean? Sum = 1/2*10*(31000+49000) sum= 5*(80,000) = 400,000

OpenStudy (anonymous):

....

OpenStudy (anonymous):

SO it's 400k + 80k?

OpenStudy (dumbcow):

where did the 80k come from?

OpenStudy (anonymous):

31+49

OpenStudy (anonymous):

So it's 400K even?

OpenStudy (anonymous):

BINGO!

OpenStudy (dumbcow):

yes sorry for the confusion

OpenStudy (anonymous):

Can you help the other Kelley? Same question, different numbers. Like 6 people are trying it now, and can't help her get this question!

OpenStudy (dumbcow):

um yeah